OpenStudy (anonymous):

1.For f(x)= 5x+1 Find(-4). a.-19 b.1 c.-21 d. 21 2. Write in Standard form an equation of the line passing through the given point with the given slope. slope=-8(-2,-2 a.8x+y=-18 b.-8x+y=-18 c. 8x-y=-18 d.8x+y=18

OpenStudy (saifoo.khan):

f(x)= 5x+1 f(-4) = 5(-4) + 1 Solve.

OpenStudy (anonymous):

a.-19 would this be the answer

jimthompson5910 (jim_thompson5910):

f(-4) = -19 is correct

OpenStudy (anonymous):

what about #2?

jimthompson5910 (jim_thompson5910):

y = mx+b -2 = -8(-2)+b -2 = 16+b -2-16 = b -18 = b b = -18 ------------- y = mx+b y = -8x-18 8x+y = -18 Answer is choice A

jimthompson5910 (jim_thompson5910):

just by coincidence, both answers are choice A

OpenStudy (anonymous):

Yeah, Thank you for the help.